the measure of angle A is 9 more than twice the measure of angle B. Ig angle A and angle B are supplementary angles, what is the measure of angle A?
Answer by nerdybill(7384)

the measure of angle A is 9 more than twice the measure of angle B. Ig angle A and angle B are supplementary angles, what is the measure of angle A?
.
Let b = measure of angle B
then
2b+9 = measure of angle A
.
b + 2b+9 = 180
3b+9 = 180
3b = 171
b = 57 degrees
.
Measure of angle A:
2b+9 = 2(57)+9 = 114+9 = 123 degrees (answer)